The simplified form of the expression
AB+ABC is:
A. AB
B. A(B+C)
C. AB+AC
1
Expert's answer
2020-12-25T15:15:48-0500

Given,

AB+ABC

We can write it as

=AB(1+c)

But we know that (1+C)=1

Hence,

AB(1+C)=AB

Hence correct answer is option A.